OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = –12 and a27 = 66.

OpenStudy (anonymous):

i got like 2.somthing

OpenStudy (anonymous):

I'm doing something wrong

OpenStudy (kira_yamato):

common difference =(66 - (-12)) / (27 - 1) = 78/26 = 3 a(42) = -12 + 3(42 - 1) = 111

OpenStudy (anonymous):

how did you get that

OpenStudy (kira_yamato):

a(n) = a(1) + (n - 1)d 66 = -12 + d(27 - 1) 78 = 26d d = 3